OpenStudy (anonymous):

Simplify the following product of complex numbers. (5 + 3i) · (6 – 2i)

OpenStudy (curry):

36+8i

OpenStudy (anonymous):

(5+3i)(6-2i) FOIL 30 - 10i + 18i + 6 36 + 8i

OpenStudy (angela210793):

=30+8i+6=36+8i